OpenStudy (anonymous):

preform the indicated operation of subtraction on the two rational expressions and reduce to lowest terms ((4x+5)/(x^2-4x+4)) - ((3)/(x-2))

OpenStudy (asnaseer):

\[\frac{4x+5}{x^2-4x+4}-\frac{3}{x-2}\]note that:\[x^2-4x+4=(x-2)^2\]can you solve this now?

OpenStudy (anonymous):

-11?

OpenStudy (asnaseer):

how did you get -11?

OpenStudy (anonymous):

from \[x+11\div(x-2)^2\]

OpenStudy (asnaseer):

the equation you just wrote down is the correct answer. but it does not simplify down to -11.

OpenStudy (asnaseer):

so answer is:\[\frac{x+11}{(x-2)^2}\]

OpenStudy (anonymous):

oh okay thanks

OpenStudy (asnaseer):

yw